# Jinhui Bai Andreas Christmann and Lei Shi

### v1

@@ 1.2 Algorithm Overview and Main Contributions
The remainder of the paper is structured as follows. In Section 2, we introduce the basic assumptions on loss functions and briefly review the theoretical background of SBFs. We then propose the T-kernel SGD with general losses. We introduce the mathematical framework underlying T-kernel SGD and present its convergence behaviors in Section 3. In Section 4, we validate the theoretical guarantees and analyze the computational complexity through numerical experiments. All proofs of the theorems are deferred to the Appendix.
@@ 2 Preliminaries and Algorithm
In this section, we outline the basic assumptions on the loss functions and give examples that satisfy them. We then review the theoretical foundations of spherical radial basis functions and their role in defining the hypothesis space. Finally, we introduce truncated-kernel stochastic gradient descent and discuss its extension to broader input domains.
@@ 2.3 Truncated Kernel Stochastic Gradient Descent
From equation (2.10), the projection operation on $\hat{g}_{n}$, i.e., $\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})$, essentially only involves operations on the coefficients of the expansion of $\hat{g}_{n}$. Therefore, in the recursion of the T-Kernel SGD (2.9), aside from computing the function value $\hat{f}_{n}(X_{n})=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{(n)}Y_{k,j}(X_{n})$, all other operations are performed on the coefficients of the basis functions $\{Y_{k,j}\}$. The explicit forms of the basis functions $\{Y_{k,j}\}$, as well as the normalization constants and related details, are provided in subsubsection A.1.1. Therefore, we can directly present the T-Kernel SGD, which works with the coefficients of the expansion, in Algorithm 1.
@@
$\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})=\left\{\begin{aligned} &\frac{Q}{\|g_{n}\|_{K}}\cdot\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}^{(n)}Y_{k,j},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}>Q,\\ &\hat{g}_{n},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}\leq Q.\end{aligned}\right$.
@@
we adopt the same parameter settings as in the original T-kernel SGD, namely $L_{n}=\min\newline \left\{k\,\Big|\,\dim{\Pi_{k}^{d}}\geq n^{\theta}\right\}$ with $0<\theta<\frac{1}{2}$ and $\gamma_{n}=\gamma_{0}n^{-t}$ for $t\in\left[\frac{1}{2},1\right)$. In addition, we employ the $\alpha$-suffix averaging scheme $\bar{f}_{\alpha n}:=\frac{1}{\alpha n}\sum_{k=(1-\alpha)n}^{n-1}\hat{f}_{k}$ as the output of the estimator.
@@ 3 Theoretical Results
$f^{*}:=\arg\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell\left(f\circ F\left(X\right),Y\right)\right],$
@@
    $\mathcal{E}(f)-\mathcal{E}(f^{*})=\mathbb{E}\left[\ell(f\circ F(X),Y)\right]-\mathbb{E}\left[\ell(f^{*}\circ F(X),Y)\right]\leq\frac{L}{2}\left\|f\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}.$ (3.1)
@@ 3.1 Optimal Rates for Excess Risk
where $\hat{f}_{n}$ denotes the last iterate in (2.11) and $\bar{f}_{\alpha n}=\frac{1}{\alpha n}\sum_{k=(1-\alpha)n}^{n-1}\hat{f}_{k}$ is the $\alpha$-suffix average. Here, $0<A_{2}\leq 1\leq A_{1}$ denote the upper and lower bounds, of $a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}$, respectively, i.e.,
@@
In online nonparametric regression, most studies on minimax optimality have focused almost exclusively on the least-squares loss, whereas investigations of general loss functionn, particularly non-convex ones, remain limited. Nevertheless, classical kernel SGD typically suffers from the saturation phenomenon, where the convergence rate ceases to improve once the regularity of the minimizer $f^{*}$ exceeds a certain threshold. For unregularized kernel SGD, [63] established convergence rates of $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\log n\right)$ for the regularity parameter $r\in(0,\frac{1}{2}]$, while [24] obtained optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\right)$ using the capacity parameter $s$, valid for $r\in\left[\frac{1}{2},1-\frac{1}{4s}\right]$. By employing Polyak averaging, [19] enhanced the robustness of the estimator and established optimal convergence rates $\mathcal{O}\left(n^{-\frac{4sr}{4sr+1}}\right)$, which depend on the capacity parameter $s$, for $r\in\left[\frac{1}{2}-\frac{1}{4s},1-\frac{1}{4s}\right]$. Incorporating an additional regularization scheme into kernel SGD helps alleviate saturation. In particular, [56] analyzed regularized kernel SGD and obtained the optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log\frac{2}{\alpha}\right)^{4}\right)$ with probability at least $1-\alpha$ for $r\in\left[\frac{1}{2},1\right]$. In contrast to previous analyses, which experience saturation when the regularity parameter $r>1$, our algorithm, when specialized to the least-squares case, effectively overcomes this phenomenon. For general loss functions, however, the nonlinear structure of the Hessian introduces substantial challenges in analyzing the convergence of kernel SGD. In online learning, classical SGD analysis yields only the slow rate $\mathcal{O}(n^{-\frac{1}{2}})$, corresponding to saturation at $r=\frac{1}{2}$. Leveraging stronger regularity conditions ($r>\frac{1}{2}$) to accelerate kernel SGD has remained an open problem. Theorem 1 shows that, under suitable regularity assumptions, T-kernel SGD attains fast rates and— to the best of our knowledge—provides the first saturation-free guarantees for online learning with general losses.
@@
In the following, we investigate the optimality of the storage complexity. The relevant definitions and concepts employed in the discussion of the lower bound on storage complexity are adapted from Section 6.3 of [64]. We now adopt a description analogous to a (probabilistic) Turing machine to formally define the general estimator. An estimator can be viewed as a mapping $G_{n}$ from the sample space $\{(X_{i},Y_{i})\}_{1\leq i\leq n}\subset\left(\mathcal{X}\times\mathcal{Y}\right)^{n}$ to the function space $f_{n}\in\mathcal{W}$. Any estimator implementable on a computer necessarily involves an encoding–decoding procedure: the encoder $E_{n}$ maps the samples $\{(X_{i},Y_{i})\}$ to a binary sequence $b_{n}$, which is stored in memory, and the decoder $D_{n}$ translates the stored $b_{n}$ into the output function $\hat{f}_{n}$. In general, as the sample size increases, the estimator produces more accurate outputs, which in turn leads to an increase in the length of the binary sequence $b_{n}$. More specifically, we introduce the following definition of the general estimator.
@@
For $l_{n}\in{\mathbb{N}}_{+}$, we define an $l_{n}$-sized estimator $G_{n}=D_{n}\circ E_{n}:(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{W}$, that is, the composition of the encoder $E_{n}$ and the decoder $D_{n}$.
@@
    For $n\in{\mathbb{N}}_{+}$, one may consider an encoding map $E_{n}:(\mathcal{X}\times\mathcal{Y})^{n}\to\{0,1\}^{l_{n}}$, which can be randomized or deterministic.
@@
Combining the above definitions, one can derive a lower bound on the complexity of the algorithm storage while achieving the minimax rate.
@@
Consider an integer sequence $\{l_{n}\}$ such that $l_{n}=o\left(n^{\frac{1}{2s(2r+1)}}\right)$ with $s>\frac{1}{2},\ r\geq\frac{1}{2}$, and let $G(l_{n})$ denote the collection of all $l_{n}$-sized estimators, one has
$\lim_{n\to\infty}\inf_{G_{n}\in G(l_{n})}\sup_{f^{*}\in\mathcal{W}\cap L_{\omega,K}^{r}(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right))}\mathbb{E}\left[n^{\frac{2r}{2r+1}}\|M_{n}\left(\{(X_{i},Y_{i})\}_{1\leq i\leq n}\right)-f^{*}\|_{\omega}^{2}\right]=\infty.$
@@ 4.1 Robust Regression on the Circle
$K_{Gaussian}(r)=\exp\left(\frac{r^{2}}{2}\right),$
$K_{Matern}^{5/2}(r)=\left(1+\sqrt{5}r+\frac{5r^{2}}{3}\right)\exp\left(-\sqrt{5}r\right).$
@@ 4.3 Binary Classification of High-Dimensional Non-Spherical Data
Figure 4: On the left is the sample–accuracy plot, and on the right is the runtime–accuracy plot.

### v3

@@ 1.2 Algorithm Overview and Main Contributions
The remainder of the paper is structured as follows. In Section 2, we introduce the basic assumptions on loss functions and briefly review the theoretical background of SBFs. We then propose the T-kernel SGD with general losses. We introduce the mathematical framework underlying T-kernel SGD and present its convergence behaviors in Section 3. In Section 4, we validate the theoretical guarantees and analyze the computational complexity through numerical experiments [^2] The code for the experimental section of the paper is located at the link: https://github.com/Researcher-Bai-1/T-kernel-SGD-with-general-losses.git. All proofs of the theorems are deferred to the Appendix.
@@ 2 Preliminaries and Algorithm
In this section, we outline the basic assumptions on the loss functions and give examples that satisfy them. We then review the theoretical foundations of spherical radial basis functions and their role in defining the hypothesis space. Finally, we introduce truncated kernel stochastic gradient descent and discuss its extension to broader input domains.
@@ 2.3 Truncated Kernel Stochastic Gradient Descent
From equation (2.10), the projection operation on $\hat{g}_{n}$, i.e., $\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})$, essentially only involves operations on the coefficients of the expansion of $\hat{g}_{n}$. In the recursion of the T-Kernel SGD (2.9), aside from computing the function value $\hat{f}_{n}(X_{n})=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{(n)}Y_{k,j}(X_{n})$, all other operations are performed on the coefficients of the basis functions $\{Y_{k,j}\}$. The explicit forms of the basis functions $\{Y_{k,j}\}$, as well as the normalization constants and related details, are provided in subsubsection A.1.1. Therefore, we can directly present the T-Kernel SGD, which works with the coefficients of the expansion, in Algorithm 1.
@@
$\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})=\left\{\begin{aligned} &\frac{Q}{\|\hat{g}_{n}\|_{K}}\cdot\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}^{(n)}Y_{k,j},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}>Q,\\ &\hat{g}_{n},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}\leq Q.\end{aligned}\right$.
@@
we adopt the same parameter settings as in the original T-kernel SGD, namely $L_{n}=\min\newline \left\{k\,\Big|\,\dim{\Pi_{k}^{d}}\geq n^{\theta}\right\}$ with $0<\theta<\frac{1}{2}$ and $\gamma_{n}=\gamma_{0}n^{-t}$ for $t\in\left[\frac{1}{2},1\right)$. In addition, we employ the $\alpha$-suffix averaging scheme $\bar{f}_{\alpha n}:=\frac{1}{\alpha n}\sum_{i=(1-\alpha)n}^{n-1}\hat{f}_{i}$ as the output of the estimator.
@@ 3 Theoretical Results
$f^{*}:=\arg\min_{f\in\mathcal{W}}\mathcal{E}(f)=\arg\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell\left(f\circ F\left(X\right),Y\right)\right],$
@@
    $\mathcal{E}(f)-\mathcal{E}(f^{*})\leq\frac{L}{2}\left\|f\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}.$ (3.1)
@@ 3.1 Optimal Rates for Excess Risk
where $\hat{f}_{n}$ denotes the last iterate in (2.11) and $\bar{f}_{\alpha n}=\frac{1}{\alpha n}\sum_{i=(1-\alpha)n}^{n-1}\hat{f}_{i}$ is the $\alpha$-suffix average. Here, $0<A_{2}\leq 1\leq A_{1}$ denote the upper and lower bounds, of $a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}$, respectively, i.e.,
@@
In online nonparametric regression, most studies on minimax optimality have focused almost exclusively on the least-squares loss, whereas investigations of general loss function, particularly non-convex ones, remain limited. Nevertheless, classical kernel SGD typically suffers from the saturation phenomenon, where the convergence rate ceases to improve once the regularity of the minimizer $f^{*}$ exceeds a certain threshold. For unregularized kernel SGD, [63] established convergence rates of $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\log n\right)$ for the regularity parameter $r\in(0,\frac{1}{2}]$, while [24] obtained optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\right)$ using the capacity parameter $s$, valid for $r\in\left[\frac{1}{2},1-\frac{1}{4s}\right]$. By employing Polyak averaging, [19] enhanced the robustness of the estimator and established optimal convergence rates $\mathcal{O}\left(n^{-\frac{4sr}{4sr+1}}\right)$, which depend on the capacity parameter $s$, for $r\in\left[\frac{1}{2}-\frac{1}{4s},1-\frac{1}{4s}\right]$. Incorporating an additional regularization scheme into kernel SGD helps alleviate saturation. In particular, [56] analyzed regularized kernel SGD and obtained the optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log\frac{2}{\alpha}\right)^{4}\right)$ with probability at least $1-\alpha$ for $r\in\left[\frac{1}{2},1\right]$. In contrast to previous analyses, which experience saturation when the regularity parameter $r>1$, our algorithm, when specialized to the least-squares case, effectively overcomes this phenomenon. For general loss functions, however, the nonlinear structure of the Hessian introduces substantial challenges in analyzing the convergence of kernel SGD. In online learning, classical SGD analysis yields only the slow rate $\mathcal{O}(n^{-\frac{1}{2}})$, corresponding to saturation at $r=\frac{1}{2}$. Leveraging stronger regularity conditions ($r>\frac{1}{2}$) to accelerate kernel SGD has remained an open problem. Theorem 1 shows that, under suitable regularity assumptions, T-kernel SGD attains fast rates and—to the best of our knowledge—provides the first saturation-free guarantees for online learning with general losses.
@@
In the following, we investigate the optimality of the storage complexity. The relevant definitions and concepts employed in the discussion of the lower bound on storage complexity are adapted from Section 6.3 of [64]. We now adopt a description analogous to a (probabilistic) Turing machine to formally define the general estimator. An estimator can be viewed as a mapping $G_{n}$ from the sample space $\{(X_{i},Y_{i})\}_{1\leq i\leq n}\subset\left(\Omega\times\mathcal{Y}\right)^{n}$ to the function space $f_{n}\in\mathcal{W}$. Any estimator implementable on a computer necessarily involves an encoding–decoding procedure: the encoder $E_{n}$ maps the samples $\{(X_{i},Y_{i})\}_{1\leq i\leq n}$ to a binary sequence $b_{n}$, which is stored in memory, and the decoder $D_{n}$ translates the stored $b_{n}$ into the output function $\hat{f}_{n}$. In general, as the sample size increases, the estimator produces more accurate outputs, which in turn leads to an increase in the length of the binary sequence $b_{n}$. More specifically, we introduce the following definition of the general estimator.
@@
For $l_{n}\in{\mathbb{N}}_{+}$, we define an $l_{n}$-sized estimator $G_{n}=D_{n}\circ E_{n}:(\Omega\times\mathcal{Y})^{n}\to\mathcal{W}$, that is, the composition of the encoder $E_{n}$ and the decoder $D_{n}$.
@@
    For $n\in{\mathbb{N}}_{+}$, one may consider an encoding map $E_{n}:(\Omega\times\mathcal{Y})^{n}\to\{0,1\}^{l_{n}}$, which can be randomized or deterministic.
@@
By combining the above definitions, one can derive a lower bound on the storage complexity while achieving the minimax rate.
@@
Consider a positive integer sequence $\{l_{n}\}$ such that $l_{n}=o\left(n^{\frac{1}{2s(2r+1)}}\right)$ with $s>\frac{1}{2},\ r\geq\frac{1}{2}$, and let $G(l_{n})$ denote the collection of all $l_{n}$-sized estimators, one has
$\lim_{n\to\infty}\inf_{G_{n}\in G(l_{n})}\sup_{f^{*}\in\mathcal{W}\cap L_{\omega,K}^{r}\left(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)\right)}\mathbb{E}\left[n^{\frac{2r}{2r+1}}\|G_{n}\left(\{(X_{i},Y_{i})\}_{1\leq i\leq n}\right)-f^{*}\|_{\omega}^{2}\right]=\infty.$
@@ 4.1 Robust Regression on the Circle
$K_{Gaussian}(r)=\exp\left(\frac{r^{2}}{2}\right),\quad K_{Matern}^{5/2}(r)=\left(1+\sqrt{5}r+\frac{5r^{2}}{3}\right)\exp\left(-\sqrt{5}r\right).$
@@ 4.3 Binary Classification of High-Dimensional Non-Spherical Data
Figure 4: The two plots above show the sample-accuracy and time-accuracy, respectively.